On January 1,2021, Krizelle Company acquired 30% of the outstanding ordinary shares of Arah Company for ₱6,000,000. This investment gave Krizelle the ability to exercise significant influence over Arah. The book value of the acquired shares was ₱5,000,000. The excess of the cost over book value was attributed to a depreciable asset which was undervalued on Arah statement of financial position and which had a remaining useful life of eight years.

For the year ended December 31,2021. Arah’ share capital outstanding is as follows:

10% cumulative preference share capital ₱3,000,000

Ordinary share capital 6,000,000

Arah reported net income of ₱2,500,000 for the year ended December 31,2021


CASE NO. 2: assuming the cumulative preference share is accounted as financial liability by Arah, answer the following:

1. What amount should krizelle record as investment income for the year ended December 31,2021?

2. What amount should krizelle record as investment in associate for the year ended December 31,2021?



1
Expert's answer
2021-06-15T16:15:34-0400

Solution:

1.). The amount that Krizelle should record as investment income for the year ended December 31, 2021, = ₱750,000

The equity method of investment calculation is applied since Krizelle company has a significant influence over Arah company by owning more than 20% interest.


Investment income = 2,500,000 "\times" 30% = 750,000

Dr. Investment in Arah’s Co. shares                              750,000

             Cr. Investment income                                                                750,000

(Being revenue made from the investment)


The amount Krizelle should record as investment income for the year ended December 31, 2021, = ₱750,000

 

2.). The amount that Krizelle should record as an investment in associate for the year ended December 31, 2021, = ₱6,615,000


Purchase price (cost) = 6,000,000

Share of net income = "2,500,000\times 30\% = 750,000"

Share of dividend received = "450,000\times 30\% = 135,000"


Dr. Investment in Arah’s Co. shares                              6,000,000

             Cr. Cash                                                                                         6,000,000

(To record the investment in Arah Co.)

Dr. Investment in Arah’s Co. shares                               750,000

             Cr. Investment income                                                                     750,000

(To record share of net income)

Dr. Cash                                                                           135,000

             Cr. Investment in Arah’s Co. shares                                               135,000

(To record dividend received)


Investment in associate = Purchase cost + share of net income – share of dividends

= 6,000,000 + 750,000 – 135,000 = ₱6,615,000